### v1

# A Large Language Model-Enabled Control Architecture for Dynamic Resource Capability Exploration in Multi-Agent Manufacturing Systems

Abstract
Manufacturing environments are becoming more complex and unpredictable due to factors such as demand variations and shorter product lifespans. This complexity requires real-time decision-making and adaptation to disruptions. Traditional control approaches highlight the need for advanced control strategies capable of overcoming unforeseen challenges, as they demonstrate limitations in responsiveness within dynamic industrial settings. Multi-agent systems address these challenges through decentralization of decision-making, enabling systems to respond dynamically to operational changes. However, current multi-agent systems encounter challenges related to real-time adaptation, context-aware decision-making, and the dynamic exploration of resource capabilities. Large language models provide the possibility to overcome these limitations through context-aware decision-making capabilities. This paper introduces a large language model-enabled control architecture for multi-agent manufacturing systems to dynamically explore resource capabilities in response to real-time disruptions. A simulation-based case study demonstrates that the proposed architecture improves system resilience and flexibility. The case study findings show improved throughput and efficient resource utilization compared to existing approaches.
I Introduction
Modern manufacturing systems face increasing complexity due to unforeseen operational disruptions (e.g. equipment failures, unexpected demand shifts) [1]. To keep competitiveness, manufacturers have been integrating advanced automation and control systems to increase adaptability [2]. However, many existing control approaches are rigid, unable to adapt to real-time variability on the shop floor, which causes bottlenecks and delays. To improve adaptability, Multi-Agent System (MAS) has been proposed to enable the coordination of products and resources [3]. By distributing decision-making across agents representing different components, such as products, machines, and robots, MAS strategies enhance flexibility in manufacturing [4].
While MASs enhance manufacturing flexibility, significant challenges exist. One challenge is the lack of real-time decision-making for dynamic reconfiguration. Manufacturing settings produce vast amounts of dynamic data from systems, products, and resources. Most systems cannot process this data in real-time for decision-making [5]. Therefore, in MAS, an adaptive control architecture is needed to respond to unanticipated disturbances (RCh1). Moreover, limited contextual awareness of MAS presents another challenge. Contextual information, such as resource locations and process parameters, is processed by predefined rules and algorithms in manufacturing [5]. These rule-based models limit the ability to flexibly adapt control strategies and interpret complex contextual changes (RCh2). Hence, MAS face challenges in improving operational efficiency based on real-time shop floor conditions. Agents also operate within certain defined ranges of capabilities determined at the initial state. This constraint limits agents utilizing wider capabilities, preventing them from adapting to changing situations [6]. Most systems require Subject Matter Expert (SME) to manually change these boundaries (RCh3). Dealing with these challenges requires a framework that allows distributed agents to learn and explore wider capabilities [6].
Figure 1: Overview of a control architecture that consists of multiple local agents including a central controller agent for manufacturing systems [6]
Large Language Models (LLMs), such as GPT-4o [7] and Llama [8], present opportunities to enhance adaptability. LLMs can support different tasks on the shop floor, enabling more effective communication. Previous studies have shown how LLMs can increase system flexibility and facilitate manufacturing interactions [9, 10]. LLMs have demonstrated the ability to interpret real-time data, support human-robot collaboration, and adapt to evolving operational requirements. Using these strengths, LLMs can be integrated into Central Controller Agent (CCA) framework for MAS, previously introduced in [6], as presented in Figure 1. This integration can improve real-time adaptability and dynamic exploration of resource capabilities.
This paper presents an LLM-enabled control architecture for MASs, enabling dynamic exploration of resource capabilities to adapt to real-time disruptions. The major contribution of this paper include: (1) designing a control architecture that enables dynamic decision-making for unexpected disruptions, (2) developing a context-aware system that evaluates the current state of the environment and performance to make informed control decisions, and (3) introducing a mechanism for dynamically exploring resource capabilities, improving agent coordination for system resilience.
The remainder of this paper is organized as follows. Section II reviews control approaches and recent applications of LLMs in manufacturing for adaptation. Section III provides an overview of MAS and defines the problem. Section IV describes the design of the control framework. Section V demonstrates the proposed framework through a case study. Finally, Section VI summarizes the paper and suggests directions for future research.
II Background
II-A Dynamic Adaptation for Manufacturing System Control
Different control strategies have been designed to enhance adaptability in manufacturing. A centralized control approach is used for resource allocation and scheduling [11, 12], leveraging sensor data for decision-making and rerouting parts during disruptions. However, as system complexity grows, centralized architectures face challenges in learning new information as they depend on predefined capabilities [6]. Distributed control strategies, such as MAS, have been introduced to tackle these challenges, where various agents make local decisions under coordination to fulfill global objectives [3, 4]. Recent MAS architectures integrate Product Agents (PAs) and Resource Agents (RAs) to enhance flexibility [13]. PAs manage product-specific decision-making, while RAs control shop-floor resources like robots and machines [13]. Existing MAS solutions improve adaptability by PAs dynamically exploring RAs for task execution [14], while RAs optimize resource allocation during disruptions [15].
Although these systems improve adaptability, they remain limited in their initial configurations. These approaches do not address extending capabilities outside of specified limits [6]. If a resource is initially configured with a subset of its possible functions, existing systems can only adapt within these predefined capabilities. Furthermore, MAS lacks a global perspective, where local adaptations are done without a holistic view. This paper introduces a CCA within MAS, leveraging an LLM to assess real-time system conditions, learn from disturbances, and dynamically extend capabilities beyond initial configurations.
II-B Large Language Models in Manufacturing for Adaptation
The introduction of LLMs has provided possibilities for enhancing adaptability across various aspects of manufacturing. Makatura et al. [16] and Li et al. [17] demonstrate how LLM translate natural language instructions into executable tasks, supporting flexible process planning and adaptive CAD/CAM workflows. By error-assisted fine-tuning and intelligent retrieval, Xia et al. [18] and Freire et al. [19] further improve domain-specific programming and knowledge sharing. Garcia et al. [20] and Fan et al. [21] use LLM for data processing and decision-making, allowing real-time task adjustments based on production constraints. Ni et al. [22] investigate LLM-based process planning, highlighting LLMs’ role in increasing adaptability across manufacturing processes.
While these studies highlight the adaptability of LLMs in manufacturing, they focus on simplified processes such as translating basic instructions, process planning, and knowledge retrieval. They do not focus on large-scale manufacturing systems, particularly in multi-agent environments where interactions between agents affect system performance. A gap exists in utilizing LLMs for context-aware decision-making and exploring resource capabilities in response to unexpected disruptions. For instance, if multiple machines operate with varying efficiency levels, LLM should not only retrieve performance data but also dynamically assess current states and constraints to enable reconfigurations. Therefore, we introduce a control architecture that utilizes LLMs for large-scale manufacturing operations, explore capabilities, and enable dynamic decision-making.
III Overview and Problem Statement
Figure 2: A Framework of Central Controller Agent for Multi-Agent Manufacturing Systems
The section provides an overview of MAS, which utilizes PAs and RAs, defined in [13, 14] and outlines the resource capability exploration problem.
III-A Overview of Multi-Agent Manufacturing Systems
The majority of MASs include PAs and RAs [13]PAs are responsible for physical parts, making decisions about processing sequences [13]. RAs manage and control physical resources, such as machines or robots, executing tasks requested by PAs [14]PAs coordinate with RAs to explore available resource capabilities and make informed decisions for guiding physical parts [14].
III-A1 Resource Agents
Each RA is equipped with a Finite State Machine (FSM) model that specifies its capabilities and interactions with neighboring RAs, which are RAs that are sharing common states for coordinated operations. The capabilities model, $M$, is formally defined for RA and contains the following elements [14]:
$M=(X,E,Tr,Prp_{p},x_{i},X_{m})$:
  • $X=\{x_{0},x_{1},\dots,x_{n}\}$ a set of states that a physical part can achieve when processed by the resource
  • $E=\{e_{0},e_{1},\dots,e_{l}\}$ is a set of events that the RA can start, each corresponding to a process in the physical system
  • $Tr:X\times E\rightarrow X$ a function that maps state-event pairs to resulting states
  • $Prp_{p}:X\rightarrow P_{p}$ a function that maps each state to physical properties that change the part
  • $x_{i}\in X$: the initial state of the part when the resources evaluate a bid request
  • $X_{m}\subseteq X$: a set of states that the part must reach to complete a task, updated with each bid request
RAs update their capability models using shop floor data and agents [14]. RAs use shared states $X_{s}\subseteq X$ to communicate, representing conditions accessible to multiple resources. Each RA maintains a table $N:X_{s}\rightarrow 2^{RA}$, which allows PAs to forward task execution requests to neighboring RAs for task allocation [14].
III-A2 Product Agents
The PA utilizes three main components to support decision-making and interaction with the RA: the process plan, product history, and environment model.
The process plan represents a customer order as an ordered list of physical property sets: $P_{d}=(P_{d1},P_{d2},\dots,P_{dn})$. Each element $p_{ij}\in P_{p}\cup P_{np}$ defines a desired property, where $P_{p}$ changes the part’s composition, and $P_{np}$ does not (e.g., location, orientation). The plan is sequential, requiring completion of $P_{1}$ before $P_{2}$, and so on.
The product history represents the current state and progression of a part. Its definition is: $PH=(X_{v},Prp_{v},RA_{c})$, where $X_{v}=\{x_{0},x_{1},\dots,x_{m}\}$ denotes the sequence of achieved states, $Prp_{v}:X_{v}\to P_{p}\cup P_{np}$ represents the set of completed physical properties, and $RA_{c}$ is the RA that informed the PA of a visited state. As the product moves through different processes, each RA updates the PA with information about state transitions and physical properties.
The environment model $M_{e}$ provides PA with a holistic view of the manufacturing environment. The environment model $M_{e}$ is defined as:
$M_{e}=(X_{e},E_{e},Tr_{e},Prp_{p,e},Ag_{h,e},x_{c})$ where $x_{c}$ is the current state of the physical part. $Ag_{h,e}$ is the event-agent association function, mapping each event to the responsible RA. The state set $X_{e}$, event set $E_{e}$, state transition function $Tr_{e}$, and physical properties $Prp_{p,e}$ follow the same definitions in the capabilities model $M$.
Additional details and examples of RA and PA models are provided in [13].
III-B Problem Statement
The dynamic capability exploration problem is the process of adapting an RA in response to disruptions by extending its operational capabilities beyond predefined boundaries. When an RA cannot fulfill its assigned tasks, the system must identify an exploration agent $RA_{e}$, that can take over the disrupted operations. This involves reconfiguring capabilities, e.g. extending a robot’s reach or adjusting a CNC spindle speed if possible or needed. To formally define this problem, we introduce the following:
$\mathcal{F}(RA_{d},X_{d},E_{d})=(RA_{e},C_{e})\rightarrow(X_{e},E_{e})$ (1)
where $RA_{d}$ is the disrupted agent unable to perform its operations, and $X_{d}$ and $E_{d}$ represent the initial states and events before the disruption. The function $\mathcal{F}$ identifies an exploration agent $RA_{e}$ and its explored capabilities $C_{e}$ (e.g. movement locations, process parameters). These capabilities define the set of feasible states, $X_{e}$, and the corresponding executable events, $E_{e}$. This formulation guarantees exploration agent, $RA_{e}$, can meet the required process plan $P_{d}$.
For the exploration agent, $RA_{e}$, to take over disrupted tasks, each state transition must follow a sequence that support process completion. These valid state transitions ensure resource progress through the states to fulfill the assigned task. This requirement is expressed as:
$\forall x_{i}\in X_{e},\forall e_{l}\in E_{e},\quad Tr(x_{i},e_{l})=x_{j}\land x_{j}\in X_{m}$ (2)
where $Tr(x_{i},e_{l})=x_{j}$ represents the state transition function, mapping a state-event pair to a new state. The condition $x_{j}\in X_{m}$ ensures that the resulting state $x_{j}$ belongs to the set of marked states, which represent states for task completion. This guarantees that new capabilities of the exploration agent $RA_{e}$ can achieve process plan $P_{d}$. To meet these requirements, this paper proposes a control architecture that enables dynamic resource capability exploration.
IV Architecture
This section presents a CCA framework for MAS, as depicted in Figure 2, which enables RAs to explore dynamic resource capabilities. The  CCA framework consists of three main components: a knowledge base, a decision-maker, and a communication manager.
IV-A Central Controller Agent Knowledge Base
The knowledge base of CCA gathers knowledge from PAs and RAs. The CCA dynamically explores resource capabilities based on real-time data by utilizing information from local agents.
IV-A1 System Knowledge
The CCA collects information from RAs regarding their status, capabilities, and performance. This includes each RA’s capabilities, current tasks indicating which product each RA is processing, and monitors resource status to identify operational disruptions. The CCA also gathers information from product history $PH$, process plan $P_{d}$, environment model $M_{e}$ to identify exploration agent $RA_{e}$. The product history $PH$ provides a record of achieved states $X_{v}$, the set of completed physical properties $Prp_{v}$, and the last resource agent $RA_{c}$ that informed the PA.
IV-A2 Performance Monitoring
By monitoring resource capacity, the CCA collects performance metrics to enhance decision-making. The CCA gathers data at discrete time steps from RAs, constructing a performance vector $PM_{i}(t)$ that includes performance indicators such as throughput, utilization, breakdown status, and availability. This is defined as $PM_{i}(t)=[m_{1}(t),m_{2}(t),\dots,m_{n}(t)]$, where each element $m_{j}(t)$ represents a specific metric measured at discrete time $t$. Continuous data updates guide the CCA’s informed decision with both current states and performance.
IV-A3 Operational Constraints
The CCA maintains operational boundaries and safety constraints that determine the permitted limits within each RA. These limits represent the thresholds for resource operation (e.g. motion range, load capacity) to guarantee feasible and safe exploration. Feasible exploration enables an RA to function within its boundaries without exceeding constraints, such as a robot arm operating within its workspace. Safe exploration prevents hazardous states, such as spindle speed or feed rate of a CNC remaining within safe limits to prevent tool breakage. The operational boundaries determine the range of allowable actions, while the safety constraints prevent hazardous behavior. The operational boundaries for a feasible exploration of a resource $i$ can be represented as ${B}_{i}=\{b_{i1},b_{i2},\dots,b_{in}\}$, where each element $b_{ij}$ corresponds to a specific operational parameter with defined minimum and maximum limits $b_{ij}\in[b_{ij}^{\min},b_{ij}^{\max}]$. Each resource is also subject to safety constraints to prevent unsafe states. The safety constraints for a safe exploration of a resource $i$ can be represented as: ${S}_{i}=\{s_{i1},s_{i2},\dots,s_{ik}\}$, with defined minimum and maximum limits $s_{ik}\in[s_{ik}^{\min},s_{ik}^{\max}]$. The feasibility of a resource executing an explored capability is determined using the validation function discussed in Section IV-B3.
IV-B Decision Maker
Using the knowledge detailed in Section IV-A, the decision maker is responsible for authorizing agents with explored capabilities. Figure 3 illustrates the LLM-driven decision-making process, which consists of three key stages: input preparation, LLM for exploration, and output generation.
IV-B1 Inputs
Figure 3: Interface for LLM-Enabled Decision-Making Process
LLMs enhance reasoning through rationale exploration, generating and evaluating multiple reasoning paths to improve problem-solving [23]. To ensure structured and reliable outputs, we employ constraint-based prompts [24] and structured Chain-of-Thought (CoT) prompts [25] to enhance the transparency and coherence of LLM decisions. Structured CoT prompts guide the LLM to explicitly outline reasoning steps, clarifying why certain RA or capabilities are selected during disruptions. The decision-maker component utilizes three types of structured inputs for LLM:
  • Instructions: Instructions guide the LLM to analyze real-time data and explore resource capabilities while considering constraints, disruptions, and performance.
  • Data: The LLM receives disrupted agent $RA_{d}$ and its capabilities, performance vectors $PM_{i}(t)$, and product data including process plan $P_{d}$, product history $PH$, and environment model $M_{e}$.
  • Template: The template structures the LLM’s output to ensure consistency so that the capabilities can be systematically processed and applied to exploration agent $RA_{e}$.
IV-B2 LLM for Resource Capability Exploration
Various studies have applied “LLM-as-a-judge” across diverse tasks and applications for scoring, ranking, and selection [26]. Although rule-based methods (e.g. selecting the nearest available robot) provide consistent and explainable decisions, they lack adaptability to unforeseen disruptions. Additionally, LLM-based decision-making strategies, such as unconstrained reasoning, lack systematic methods to ensure consistency and reproducibility, which are critical in industrial applications. In this work, we leverage LLMs to evaluate RAs for capability exploration using a weighted scoring function:
Algorithm 1 Exploration Resource Agent Validation
1:procedure validate($O=(RA_{e},C_{e})$, $exploration\_params$, $n$)
2:  $i\leftarrow 0$, $valid\leftarrow false$, $feedback\leftarrow\text{None}$
3:  while $i<n$ and not $valid$ do
4:   $i\leftarrow i+1$
5:   if syntaxCheck(O) = false then
[…]
7:     continue
8:   $RA_{e}\leftarrow getInfo(O)$
9:   if $RA_{e}\notin model$ then
[…]
11:     continue
12:   $C_{e}\leftarrow RA_{e}.capabilities$
13:   $(b_{ij}^{\min},b_{ij}^{\max})\leftarrow exploration\_params.operation\_bounds$
14:   $(s_{ik}^{\min},s_{ik}^{\max})\leftarrow exploration\_params.safety\_limit$
[…]
16:     $valid\leftarrow true$
17:   else
[…]
19:  return $valid,feedback$
$S_{i}=\sum_{j=1}^{n}\omega_{j}\cdot f_{j}(x_{ij})$ (3)
where score $S_{i}$ represents the suitability of a candidate resource agent $i$ for exploration based on multiple performance indicators. $\omega_{j}$ is the weight assigned to factor $j$ by the LLM, and $x_{ij}$ is the raw value, such as throughput, utilization, or breakdown rates. The function $f_{j}$ normalizes these values. The agent with the highest score is selected as exploration agent $RA_{e}$, and its capabilities $C_{e}$ are adjusted by exploring the disrupted agent’s functional responsibilities. The scoring approach provides structured, quantifiable comparisons among candidates, aiding transparency and validation. Future work will investigate refinements to prompt design and methodologies to improve decision reliability.
IV-B3 Outputs
The output $O$ consists of the exploration agent $RA_{e}$ and explored capabilities $C_{e}$, which is represented as $O=(RA_{e},C_{e})$. The CCA processes this output to reconfigure the resource’s capabilities model $M$ with states $X$ and events $E$.
Although LLMs are effective for analyzing and generating information, their outputs can be inconsistent. Several studies, such as Self-Refine [27], have employed iterative refinement to reduce errors and enhance consistency. We apply a similar approach, using feedback-driven adjustments to ensure LLM decisions remain within predefined operational and safety constraints. This method is critical in manufacturing contexts, where safety constraints and operational parameters are essential to maintain system stability and reliability.
The validation process, as demonstrated in Algorithm 1, ensures that the selected agent and capabilities comply with both system constraints and safety limits. The validation begins with the CCA retrieving the operational boundary set $\mathcal{B}_{i}$ and safety constraint set $\mathcal{S}_{i}$. Before proceeding, the generated output $O$ is checked for syntax error, ensuring the output $O$ is generated based on the expected structure. If an error is detected, feedback is generated and used to refine the next iteration. The CCA then initializes the validation loop, setting the iteration counter to zero and marking the output as invalid. For each iteration, the CCA extracts information of exploration agent $RA_{e}$ from $O=(RA_{e},C_{e})$. If $RA_{e}$ is not found in the system model, the validation process continues with updated feedback. The corresponding explored capabilities, $C_{e}$ of $RA_{e}$ are then retrieved and compared against $\mathcal{B}_{i}$ and $\mathcal{S}_{i}$. If these conditions hold, the validation is successful, and the output is applied to the system. Otherwise, feedback is provided, and the process continues until a valid response is obtained or the iteration limit $n$ is reached.
Figure 4: Sequence Diagram for Exploration Process
IV-C Communication Manager
The Communication Manager is responsible for updating the capabilities model $M$ based on the output $O=(RA_{e},C_{e})$ received from the LLM. This process begins by extracting the original event set $E^{\text{original}}$ and $X_{s}$ from the selected $RA_{e}$. Disrupted events from $RA_{d}$ are removed to prevent invalid transitions.
The newly assigned $C_{e}$ are converted into $e_{j}\in E$, each defining specific state transitions. The event set $E^{\prime}$ is constructed by merging the original events with the newly assigned events, represented as:
$E^{\prime}=E^{\text{original}}\cup E^{\text{new}}$ (4)
This integration enhances the RA ’s operational efficiency by enabling more direct and flexible state transitions, as captured by the updated state transition function $Tr$, which maps initial states $x_{i}\in X$, to target states within the marked state $X_{m}$.
Additionally, neighboring agents are added to the neighbor table $N:X_{s}\rightarrow 2^{RA}$ for enhanced coordination. The PAs then cooperates with RAs through a bidding process to select the most efficient path while considering constraints and processing time [14].
IV-D Exploration Process Workflow
The sequence diagram, as depicted in Figure 4, illustrates the exploration process. The process is divided into three primary stages: Data Collection, Capability Exploration, and Capability Update.
The Data Collection stage begins with CCA detecting a disruption and sends a request for handling the event to $RA_{d}$ (Step 1a). The CCA gathers real-time information on the disruption and the resource’s performance (Step 1b). This step is performed for all affected agents within a loop.
In the Capability Exploration stage, the CCA sends a prompt to the LLM to handle disruption (Step 2a). The LLM then selects an exploration agent $RA_{e}$ and explored capabilities $C_{e}$ (Step 2b). The exploration agent $RA_{e}$ and explored capabilities $C_{e}$ undergo validation, where each capability is checked against predefined operational and safety constraints. If a capability does not meet the criteria, the agent is invalid (Step 2c), and the exploration loop continues. If all constraints are satisfied, the explored capabilities $RA_{e}$ handle disruption (Step 2d).
During the Capability Update, the CCA removes disrupted capabilities from the capability list (Step 3a) and triggers the repair process (Step 3b). The explored capabilities $C_{e}$ are added to the exploration agent $RA_{e}$ (Step 3c), expanding its functionality. The exploration agent $RA_{e}$ then proceeds to execute its tasks using the updated capabilities (Step 3d) by collaborating with PAs.
Figure 5: (a) An overview of a system with three robots, two machines, and three buffers. A disruption in Robot 2 affects Product A operations. (b) The CCA framework responds to the disruption at $X3$ and $X4$ by authorizing Robot 1 to expand its capabilities to take over the disrupted operations, ensuring task execution at $X8$.
IV-E Exploration Example
Figure 6: Case Study Setup
TABLE I: Machine Process Times
Process P1 P2 P3 P4 P5 P6
Time (Ticks) 150 60 110 100 170 20
TABLE II: Breakdown Start Times and Recovery Times
Robot Initial Breakdown (Tick) MTTR (Ticks)
M12 1000 450
B3 2500 450
B4 3000 340
B6 4500 390
Figure 5 illustrates an example of dynamic resource capability exploration in MAS with three robots (Robot 1, Robot 2, Robot 3), two machines (Machine 1, Machine 2), and three buffers (Inventory, Product A Buffer, Product B Buffer). A disruption occurs at Robot 2, preventing it from transferring Product A to Machine 1, affecting states $X3$ (at Product A Buffer) and $X4$ (moving to Machine 1). The CCA collects RAs status and performance $PM_{i}(t)$ by retrieving data from RA through direct queries. Metrics, including availability ($A_{i}$), utilization ($U_{i}$), and proximity ($P_{i}$), are stored in a knowledge base for LLM to evaluate potential exploration agents $RA_{e}$. The LLM calculates a score $S_{i}$ using a weighted function $S_{i}=w_{1}A_{i}+w_{2}(1-P_{i})+w_{3}U_{i}$ to select Robot 1 as the exploration agent $RA_{e}$ over Robot 3, based on its higher score. The CCA then validates operational boundaries by checking movement range constraints, confirming that Robot 1 can execute the required operations. The CCA assigns Robot 1 as the $RA_{e}$ and updates its capability model by integrating the disrupted events $\sigma_{3}$ and $\sigma_{4}$. The event set $E^{\prime}$ is modified, and shared states $X_{s}$ are updated, enabling Robot 1 to transition through $X8$ (Inventory to Machine 1) and reach $X5$ (at Machine 1). Through the bidding process, the PA identifies Robot 1 and state $X8$ as the most efficient path to achieve the marked state $X7$ (processed at Machine 1).
V Case Study
To evaluate the CCA architecture, an MAS was setup in a simulated environment. This section outlines the setup and showcases the exploration process in a scaled-up case study, followed by an analysis of the results.
V-A Case Study Setup
The Repast Symphony (RepastS) [28] platform is used to model and simulate MAS within the CCA architecture. The platform updates at each time step, referred to as a tick [28]. This study modifies and extends previous works [13, 29]. The simulated facility, as shown in Figure 6, consists of 20 stations (diffusion, ion implementation, and lithography) connected by a network of 9 material handling robots, each designated to its own cell or transfer stations. For example, robot B1 can only transfer parts within Cell 1, while robot M12 is responsible for transferring parts between Cell 1 and Cell 2. Buffers are positioned between the robots and also placed at each cell. Wafer lots enter the facility through an entrance and exit after completing a defined set of production steps. The facility supports six manufacturing processes (P1–P6), each performed at designated stations with processing times, as detailed in Table I. In this simulation, four robots experience disruptions starting with robot M12 at tick 1000. The initial breakdown and recovery times, including their Mean Time To Repair (MTTR), are detailed in Table II. The CCA handles disruptions as outlined in Section IV. In this case study, we employed GPT-4o by OpenAI [7].
V-B Case Study and Results
Figure 7: Throughput comparison between scenario S1 and three trials of scenario S2
Figure 8: Utilization rate comparison between the exploration agent and the disrupted agent during breakdown in S2-3
A batch of 25 parts enters the system at tick 10, following the process sequence P1 $\to$ P2 $\to$ P3 $\to$ P4 $\to$ P5 $\to$ P6. This study compares two scenarios: (S1) the existing framework of Kovalenko et al. [29], and (S2) the CCA architecture, which was tested across three trials to mitigate disruptions. The performance was evaluated by comparing the number of completed parts over time and the utilization rates of $RA_{d}$ and $RA_{e}$. The following assumptions are made for the case study:
  • A.1
    The communication delay from LLM is not considered.
  • A.2
    No two robots breaking down simultaneously.
  • A.3
    Once restored, robots revert to their original capabilities.
A.1 ensures that decision-making performance is evaluated independently without interference from variable response times. A.2 prevents overlapping disruptions, allowing a clear evaluation of the system’s response to individual failures. Finally, A.3 isolates the effects of dynamic capability exploration without additional variability.
The throughput results, shown in Figure 7, demonstrate a significant improvement with the CCA architecture in handling breakdowns, as indicated by the gray-shaded regions representing breakdown periods detailed in Table II. The existing framework (S1) completed only 15 parts, as disruptions prevented parts from finding feasible paths, leading to their removal. In contrast, the CCA architecture, tested in three trials S2-1, S2-2, and S2-3, mitigated failures, completing 23, 23, and 21 parts, respectively. Some parts still failed due to real-time disruptions, particularly when pre-planned routes became invalid or when parts were stuck in robots that had broken down, forcing their removal.
To further analyze the system behavior, the utilization rates of disrupted agent $RA_{d}$ and exploration agent $RA_{e}$ were compared, as observed in Figure 8 for S2-3, where robot M12 and robot B3 experienced breakdowns. The CCA explored new capabilities for robot B2 and robot B1, respectively. The gray-shaded regions in Figure 8 indicate the periods during which the robots experienced disruptions. During these periods, both the disrupted agents $RA_{d}$ and the exploration agents $RA_{e}$ exhibit similar behavior. $RA_{d}$ showed a decline in utilization and $RA_{e}$ demonstrated an increase as they were assigned new tasks.
V-C Insights from Case Study
The case study highlights that the proposed CCA architecture enhances system resilience and adaptability. This research demonstrated how the system can maintain efficient operations despite disruptions, minimizing part removals and reducing failures. Even when agents experience breakdowns. the system sustains high throughput and minimizes bottlenecks. These findings suggest that the CCA architecture offers a robust and flexible approach for handling dynamic challenges.
However, there are several challenges identified in this study. A major issue is the limited explainability of LLM-driven decision-making. Although agent exploration was based on a weighted scoring function, the LLM did not always make the best decisions. Future research should focus on increasing prompt effectiveness and enhancing the explainability of LLMs. For instance, the LLM occasionally assigned inconsistent weights to performance metrics, such as prioritizing utilization rate over proximity without clear justification. Therefore, reliable decision-making depends on refining the prompt design and increasing the transparency of LLM reasoning.
Another challenge is the generalizability of the proposed architecture. Although the system successfully handled unexpected robot breakdowns, further evaluation is needed to assess its applicability to other uncertainties, such as machine failures, demand shifts, and simultaneous disruptions of multiple agents. Although fine-tuning LLM can enhance accuracy in manufacturing [18] for specific scenarios, developing an approach capable of handling any disturbances is important for maintaining efficiency. Therefore, the architecture should utilize a knowledge base that is structured for MAS runtime data [30]. This can ensure the system retrieves relevant context related to real-time disturbances with LLM using Retrieval Augmented Generation (RAG) [31].
VI Conclusion and Future Work
This paper introduced a CCA architecture, achieving real-time adaptability (RCh1), enabling context-aware decision-making based on real-time data (RCh2), and relaxing pre-defined operational boundaries by dynamically exploring resource capabilities (RCh3). The proposed framework enhanced real-time adaptation and resource flexibility using LLM. The case study demonstrated that the CCA architecture effectively mitigates disruptions, improving system resilience and operational flexibility. The case study results showed the architecture enabled the system to reassign tasks to exploration agents, minimizing the impact of disruptions on production flow and throughput.
Future work will focus on enhancing the transparency and interpretability of LLM-based decisions to improve system validation and trust in industrial settings. Additional studies will also involve evaluating the architecture against more varied and complex disruptions to evaluate resilience and effectiveness compared to conventional rule-based adaptation methods. The goal is to demonstrate reliable and efficient performance in highly dynamic manufacturing scenarios.